OpenStudy (anonymous):

Geometric series; Question; 1. Summation of 5(7/8)^n where n =0

OpenStudy (sasogeek):

any number with the exponent 0 is 1, that is \(\large x^0=1 \) that means \(\large 5(\frac{7}{8})^0=5 \times 1=5 \) sum of 5 equals 5....

OpenStudy (sasogeek):

I hope there's some sense in it :)

OpenStudy (bahrom7893):

that's right

OpenStudy (sasogeek):

phew! :)